OpenStudy (anonymous):

The Problem Was -2m + 13 + 2m - 3 And I Worked It Out And Got M = 0

OpenStudy (lalaly):

no equal sign?

OpenStudy (anonymous):

Oh Yeah Heres The Problem -2m + 13 = 2m + 3

OpenStudy (lalaly):

-2m-2m=3-13 -4m=-10 m=10/4